# Emergent QED3 at the bosonic Laughlin state to superfluid transition

### v1

@@ Abstract
Here we focus on continuous phase transitions that connect a topologically ordered phase to a symmetry breaking phase. A simple pathway is anyon condensation Burnell (2018): when a bosonic anyon that braids non-trivially with all other anyons and carries a microscopic symmetry charge condenses, its condensation simultaneously confines the remaining anyons and breaks the symmetry Chubukov et al. (1994); Jalabert and Sachdev (1991); Senthil and Fisher (2000). A recent example is the transition from the bilayer $\nu=1/3+2/3$ fractional quantum Hall state to an exciton superfluid. It is driven by condensation of the fractional exciton $(e/3,-e/3)$ and therefore belongs to the XY universality class Zhang et al. (2023).
@@
Previous numerical studies, largely on lattice hopping models Barkeshli et al. (2015); Kamal et al. (2022), have not settled the issue. The transition has alternately appeared first-order, indirect, or inaccessible. In this work, we revisit the problem in a continuum setting: half-filled bosons confined to the lowest Landau level and perturbed by a lattice potential. For a weak potential the ground state is the bosonic $\nu=1/2$ Laughlin fluid; for strong potential kinetic energy dominates and a superfluid emerges despite the magnetic field. Using infinite-cylinder density-matrix renormalisation group (DMRG) we find a single, continuous Laughlin–to–superfluid transition.
@@
$H=\frac{1}{2}\int_{\mathbf{r},\mathbf{r}^{\prime}}\hat{\rho}_{\mathbf{r}}\delta_{\mathbf{r}-\mathbf{r}^{\prime}}\hat{\rho}_{\mathbf{r}^{\prime}}+4\pi V_{m}\sum_{j=1}^{2}\int_{\mathbf{r}}\hat{\rho}_{\mathbf{r}}\cos\bigl(\mathbf{G}_{j}\cdot\mathbf{r}\bigr{missing}),$ (1)
@@
$\varepsilon_{\mathbf{k}}=-4\pi e^{-\pi/2}V_{m}\sum_{j=1}^{2}\cos\bigl(\mathbf{k}\cdot\mathbf{a}_{j}\bigr{missing}),$ (2)
@@
Figure 2: Momentum–resolved correlation lengths $\xi(\mathbf{k})$ (color scale) on a cylinder of circumference $L_{y}=8a$ at $\chi=10000$. Panels (a)–(c) show the charge-neutral sector at $V_{m}=0.022$ (Laughlin), $0.084$ (critical), and $0.10$ (superfluid). Panel (d) shows the charge-$1$ sector at $V_{m}=0.10$. At criticality the correlation lengths at $X$, $Y$, and $M$ become almost equal, indicating an emergent $\mathsf{SO(3)}$ symmetry.
@@
If both partons occupy $\nu=1$ IQHE bands, the composite boson $b=cf$ realizes the $\nu=1/2$ Laughlin state. Integrating out the $c$ parton turns $f$ into the composite fermion in the standard composite–fermion picture of bosonic Laughlin state. If $f$ instead enters a $C=-1$ band while $c$ stays in $C=+1$, the Chern–Simons terms cancel, $a_{\mu}$ becomes gapless, and the boson condenses into a superfluid. We model $f$ with a square-lattice $\pi$-flux Hofstadter Hamiltonian that hosts four Dirac cones at (±$(\pm\pi/2a,\pm\pi/2a)$ as shown in Fig. 3(a). Mass terms $m_{I}$ that gap these cones (equivalent to tuning a small diagonal hopping) set the Chern number of the lower band. Integrating out $c$ yields the two-flavour QED3–Chern–Simons theory Barkeshli and McGreevy (2014); Song et al. (2024)
@@
Transitions described by the QED3–CS theory lie beyond Landau’s paradigm and are now being explored in moiré platforms including twisted homobilayer transition-metal dichalcogenides and rhombohedral pentalayer graphene. The flux-threading spectroscopy and momentum-resolved diagnostics introduced here can be applied directly to those materials.
Note added— During writing of this manuscript, we become aware of Ref. Lu et al., 2024a, b; Lotrivc and Simon, 2025, which studies a similar transition in lattice systems.
Acknowledgments. We thank Yin-Chen He, Maissam Barkeshli, and Zi-Yang Meng for useful discussions. M.Z. is supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC02-05-CH11231 (Theory of Materials program KC2301). T.W. is supported by the Heising-Simons Foundation, the Simons Foundation, and NSF grant No. PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). X.Y.S. is supported by the Gordon and Betty Moore Foundation EPiQS Initiative through Grant No. GBMF8684 at the Massachusetts Institute of Technology, and by Early Career Scheme of Hong Kong Research Grant Council with grant No. 26309524. T.S. is supported by NSF grant DMR-2206305. TS and MZ are also partially supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Grant No. 651446, TS, and Grant No. 1151944, MZ). TW is suppored by the Simons Collaboration on Ultra-Quantum Matter via MZ. This research uses the Lawrencium computational cluster provided by the Lawrence Berkeley National Laboratory (Supported by the U.S. Department of Energy, Office of Basic Energy Sciences under Contract No. DE-AC02-05-CH11231).
@@
$\varepsilon(\mathbf{k})=-2V_{m}e^{-\mathbf{G}^{2}\ell_{B}^{2}/4}\sum_{j}\cos\bigl(\mathbf{k}\!\cdot\!\mathbf{a}_{j}\bigr{missing}).$ (S14)

### v2

@@ Abstract
Here we focus on continuous phase transitions that connect a topologically ordered phase to a symmetry breaking phase. A simple pathway is anyon condensation Burnell (2018): when a bosonic anyon that braids non-trivially with all other anyons and carries a microscopic symmetry charge condenses, its condensation simultaneously confines the remaining anyons and breaks the symmetry Chubukov et al. (1994); Jalabert and Sachdev (1991); Senthil and Fisher (2000). A recent example is the transition from the bilayer $\nu=1/3+2/3$ fractional quantum Hall state to an exciton superfluid. It is driven by condensation of the fractional exciton (e/3,e/3)(e/3,-e/3) and therefore belongs to the XY universality class Zhang et al. (2023).
@@
Previous numerical studies, largely on lattice hopping models Barkeshli et al. (2015); Kamal et al. (2022); Motruk and Pollmann (2017); Zeng (2021), have not settled the issue. The transition has alternately appeared first-order, indirect, or inaccessible. In this work, we revisit the problem in a continuum setting: half-filled bosons confined to the lowest Landau level and perturbed by a lattice potential. For a weak potential the ground state is the bosonic $\nu=1/2$ Laughlin fluid; for strong potential kinetic energy dominates and a superfluid emerges despite the magnetic field. Using infinite-cylinder density-matrix renormalisation group (DMRG) we find a single, continuous Laughlin–to–superfluid transition.
@@
[…]
@@
[…]
@@
Figure 2: Momentum–resolved correlation lengths $\xi(\mathbf{k})$ (color scale) on a cylinder of circumference $L_{y}=8a$ at $\chi=10000$. Panels (a)–(c) show the charge-neutral sector at $V_{m}=0.022$ (Laughlin), $0.084$ (critical), and $0.10$ (superfluid). Panel (d) shows the charge-11 sector at $V_{m}=0.10$. At criticality the correlation lengths at $X$, $Y$, and $M$ become almost equal, indicating an emergent $\mathsf{SO(3)}$ symmetry.
@@
If both partons occupy $\nu=1$ IQHE bands, the composite boson $b=cf$ realizes the $\nu=1/2$ Laughlin state. Integrating out the $c$ parton turns $f$ into the composite fermion in the standard composite–fermion picture of bosonic Laughlin state. If $f$ instead enters a C=1C=-1 band while $c$ stays in $C=+1$, the Chern–Simons terms cancel, $a_{\mu}$ becomes gapless, and the boson condenses into a superfluid. We model $f$ with a square-lattice $\pi$-flux Hofstadter Hamiltonian that hosts four Dirac cones at $(\pm\pi/2a,\pm\pi/2a)$ as shown in Fig. 3(a). Mass terms $m_{I}$ that gap these cones (equivalent to tuning a small diagonal hopping) set the Chern number of the lower band. Integrating out $c$ yields the two-flavour QED3–Chern–Simons theory Barkeshli and McGreevy (2014); Song et al. (2024)
@@
Transitions described by the QED3–CS theory lie beyond Landau’s paradigm and are now being explored in moiré platforms including twisted homobilayer transition-metal dichalcogenides and rhombohedral pentalayer graphene. In Supplementary Material we discussed a similar bosonic FQH-SF transition at $\nu=3/4$ and relevance to moiré systems.The flux-threading spectroscopy and momentum-resolved diagnostics introduced here can be applied directly to those materials.
Note added— During writing of this manuscript, we become aware of Ref. Lu et al., 2024a, b; Lotrivc and Simon, 2025, which studies a similar transition in lattice hopping systems.
Acknowledgments. We thank Yin-Chen He, Maissam Barkeshli, and Zi-Yang Meng for useful discussions. This work is partially supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (Grant No. 651446, TS, and Grant No. 1151944, MZ). T.W. is supported by the Simons Collaboration on Ultra-Quantum Matter via M.Z., and the Heising-Simons Foundation, the Simons Foundation, and NSF grant No. PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). T.S. is also supported by NSF grant DMR-2206305. M.Z. is also supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Materials Sciences and Engineering Division under Contract No. DE-AC02-05-CH11231 (Theory of Materials program KC2301). X.Y.S. is supported by the Gordon and Betty Moore Foundation EPiQS Initiative through Grant No. GBMF8684 at the Massachusetts Institute of Technology, and by Early Career Scheme of Hong Kong Research Grant Council with grant No. 26309524. This research uses the Lawrencium computational cluster provided by the Lawrence Berkeley National Laboratory (Supported by the U.S. Department of Energy, Office of Basic Energy Sciences under Contract No. DE-AC02-05-CH11231).
@@
[…]
@@
SVII Bosonic Fractional quantum Hall(FQH) to superfluid transition at $\nu=-3/4$ and connection to Moire systems
In this section we discuss a FQH-SF transition theory at filling $\nu=3/4$ closely related to the transition studied in the main text. This theory is also connected to the putative continuous transition from composite Fermi liquid(CFL) to Fermi liquid(FL) at filling $\nu=-3/4$, potentially realizable in Moire TMD bilayer systems, at filling $\nu=-3/4$.
We start from a parton decomposition for the electrons $c_{s}=bf_{s}$ where $b,f_{s}$ are charged holons and spinons, respectively and a $U(1)$ gauge field $a$ that couples oppositely to $b,f_{s}$.This construction comes with the gauge constraint on the particle numbers $n_{c}=n_{b}=n_{f}$. To describe a composite Fermi liquid of electrons at filling $\nu=-3/4$, ref Song et al. (2024); Dong et al. (2023) proposed to think of the system as a composite of a $\nu=-1$ IQH state and $\nu=1/4$ composite Fermi liquid state. The transition to neighboring Fermi liquid is described by the transition of the $\nu=1/4$ system, from composite Fermi liquid to Fermi liquid. The resulting Fermi liquid hence has a large Hall conductivity from the $\nu=-1$ IQH sector. This is controlled by the holon transition from FQH at $\nu=1/4$ to superfluid, with the spinons filling a Fermi surface, e.g.
$\mathcal{L}=\mathcal{L}_{FQH-SF}[b;a]+\mathcal{L}_{f_{s};a},$
$\mathcal{L}_{f_{s};a}=\sum_{s}f_{s}^{\dagger}(i\partial_{t}-\mu)f_{s}+\sum_{\mu=x,y}\frac{|(i\partial_{\mu}-a_{\mu})f_{s}|^{2}}{2m},$ (S23)
where $\mathcal{L}_{FQH-SF}[b;a]$ describes the holon sector transition from ref Song et al. (2024) and $\mathcal{L}_{f_{s};a}$ describes the spinon Fermi surface state.
Here we discuss a different route: We consider the holon FQH state at $\nu=3/4$ and spinons fill a Fermi sea at $\nu=3/4$, the physical system of which is a particle-hole conjugate of $\nu=-3/4$ composite Fermi liquid as relevant to twisted MoTe2. To describe the holon FQH state, we further preform a parton for the holon $b=f_{1}f_{2}$ where $f_{1,2}$ are fermionic particles, both at the same filling $3/4$ as required by gauge constraint. $f_{1,2}$ each fill Chern bands with total Chern number $C_{1}=1,C_{2}=3$ which constitutes the bosonic Jain state of $\nu=p/(p+1)$ for $p=3$. When the Chern band for $f_{2}$ goes through a transition to Chern number $C_{2}=-1$, the bosons form a superfluid. Hence the system enters the Fermi liquid phase. Note that the Chern bands can be formed at mean-field level when $f_{1,2}$ each sees $\pm 1/4$ flux quantum per unit cell, hence quadrupling the unit cell and the corresponding effective filling is $3$. This allows the $f_{i}$’s to completely fill Chern bands.
The critical theory is controlled by the $\nu=3/4$ boson FQH to superfluid transition,
$\mathcal{L}_{b;a,\alpha}=\sum_{i=1}^{4}\bar{\psi}_{i}i\not{\partial}_{\alpha}\psi_{i}+m\sum_{i}\bar{\psi}_{i}\psi_{i}+\frac{2}{4\pi}\alpha d\alpha+\frac{1}{4\pi}ada-\frac{1}{2\pi}ad\alpha,$ (S24)
where $\alpha$ is the gauge field that couples oppositely to $f_{1,2}$ and the four Dirac fermions occur at the Chern number changing transition. When $m<0$, integrating out the Dirac fermions gives the term $\frac{2}{4\pi}\alpha d\alpha$, which upon further integrating out $\alpha$, gives $\frac{3}{16\pi}ada$, consistent with $\nu=3/4$ FQH response for the holons that couple to $a$. When $m>0$,integrating out the Dirac fermions gives the term $\frac{2}{4\pi}\alpha d\alpha$, which cancels the self Chern-Simons term for $\alpha$ and Higgs $a$. This indicates that bosons form a superfluid. The degeneracy of $4$ Dirac fermions at the transition is protected by the $1/4$ “magnetic field” flux felt by $f_{i}$’s and essentially the filling. Whether this transition respects $C_{6}$ rotation remains to be investigated.
Combining the action for the spinon, we have for the complete critical theory for a transition from composite Fermi liquid at $\nu=3/4$ and Fermi liquid,
$\mathcal{L}_{CFL-FL}=\mathcal{L}_{b;a,\alpha}+\mathcal{L}_{f_{s};a}.$
Hence a generalization of the transition studied in this report to $\nu=3/4$ will be indicative of the putative composite Fermi liquid to Fermi liquid transition in moire TMD platforms.